On December 31, 2017, Dow Steel Corporation had 650,000 shares of common stock and 35,000 shares of 9%, noncumulative, nonconvertible preferred stock issued and outstanding. Dow issued a 5% common stock dividend on May 15 and paid cash dividends of $450,000 and $74,000 to common and preferred shareholders, respectively, on December 15, 2018.

On February 28, 2018, Dow sold 64,000 common shares. Also, as a part of a 2017 agreement for the acquisition of Merrill Cable Company, another 26,000 shares (already adjusted for the stock dividend) are to be issued to former Merrill shareholders on December 31, 2019, if Merrill's 2019 net income is at least $550,000. In 2018, Merrill's net income was $680,000.

In keeping with its long-term share repurchase plan, 7,000 shares were retired on July 1. Dow's net income for the year ended December 31, 2018, was $2,350,000. The income tax rate is 40%.

As part of an incentive compensation plan, Dow granted incentive stock options to division managers at December 31 of the current and each of the previous two years. Each option permits its holder to buy one share of common stock at an exercise price equal to market value at the date of grant and can be exercised one year from that date. Information concerning the number of options granted and common share prices follows:
Problem 19-13 EPS; nonconvertible preferred stock; treasury shares; shares sold; stock dividend; options, convertible bonds;Earnings per Numerator Denominator = share Dows 2,276 735 - $ basic Dows 739 2,276 - $ diluted 3.08 *Red text indicates no
Answer #1

Numerator

/

Denominator

=

Earnings per share

Dow’s Diluted

2330

/

792

=

$2.94

Numerator = net income – preferred dividends + after-tax interest savings = 2350-74+(90*(1-40%)) = $2330

Denominator = weighted average number of shares =shares at Jan. 1 + new shares – shares retired + exercise of options + contingent shares + conversion of bonds = (650*12/12*1.05)+(64*10/12*1.05)-(7*6/12)+4+26+27 = 792

Shares

12000

Exercise price

16

$192000

average market price

$24

Shares

8000

Exercise of options (12000-8000)

4000 shares

contingent shares

conversion of bonds = 900000/1000*30 = 27000

Diluted EPS = 2330/792 = $2.94